13 coins worth $1.05.If the coins are just nickels and dimes.how many nickels are there
Answer by swincher4391(1107)

n+d = 13
.05n + .10d = 1.05
n = 13 -d

.05(13-d) + .10d = 1.05
.65 - .05d +.10d = 1.05
.05d = .40
d = 8
n = 5